OpenStudy (anonymous):

Will give medal. Simplify the expression the quantity thirty-seven minus twenty-nine, end quantity, divided by two times four. Show your work. Then, show how the solution is affected if multiplication and division are not applied from left to right.

OpenStudy (anonymous):

PLEASE HELP

OpenStudy (anonymous):

37-29/ 2*4= 8/8=1 since we have to solve the quantity first, the new problems is 8/2*4 since 37-29=8. now, 2*4=8, so 8/8=1, so new simplified expression is 1.

OpenStudy (anonymous):

THANK YOU THANK YOU THANK YUO